OpenStudy (anonymous):

An isosceles trapezoid has bases of 38.14 m and 14.32. Each base angle measures 48.13.. Find the length of each of the equal sides

OpenStudy (anonymous):

This question would be easier if you draw a diagram. draw a isos. trapezoid. Now, in it, make a rectangle. Something like this: _______ /| |\ / | | \ <----y /__|_______|__\ ^---x since the trapezoid is isosceles, we can assume that the bases are "centered" so (base1 - base2)/2 would be the length of the base of one of the triangles, let's call that x. So now you can just trignometry to find the side. (let's call the "equal sides" y") cos(48.13) = x/y y = cos(48.13)/x